Various repair and maintenance work was needed on the weather station. The rain sensor was broken by failing ice. A new sensor was ordered and in the mean time we are using an older spare that is less sophisticated and more sensitive to false alarms. The mains power supply cable from the telescope was severely affected by a lightning storm in February and as a quick fix power is now supplied through a simple cable from the service building. Taking the opportunity, a cable duct will be made from the service building where a new cable will be installed with new lightning protection. During the storm also the communication was affected as a RS232 to fibre converter was damaged and had to be replaced. There were also some problems with the fuse of the weather station cause by the new building drive system (see below).